# Bayesian evidence for two peaks in the sound speed in cold dense QCD

### v1

@@ B.5 Summary of models
This appendix examines assumptions underlie the main results in addition to those discussed in appendix B. I begin by discussing choices of prior distributions. In the main text, neutral priors without preferences for the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ are chosen. They are obtained by imposing separate flat distributions over regions where $\Delta P-\Delta P_{\mathrm{mean,max}}>0$ and where $\Delta P-\Delta P_{\mathrm{mean,max}}<0$. Normalizations of the two segments are obtained from demanding the integrated probability equals on both sides. The effect of eq. 7 is minor since a considerable fraction of samples satisfy $C_{s,\mathrm{max}}^{\mathrm{NS}}>1/3$ and $C_{s,\mathrm{min}}^{\mathrm{UD}}<1/3$ in the posterior. I have verified that neutral priors on $\Delta P-\Delta P_{\mathrm{mean,max}}$ alone lead to unbiased priors for the two-peak scenario.
Another common choice in the literature is a simple flat prior over possible ranges of $\Delta P-\Delta P_{\mathrm{mean,max}}$. It is deemed the least informative and generally preferred. Bayes factors based on this choice are shown in table 4. Here, the priors display clear preferences. The two-peak scenario is generally disfavored by these priors unless $X\gtrsim 3$. Yet the astrophysical evidence remains, albeit with less statistical significance in these disadvantageous settings.
@@
I emphasize that the phase space for the two-peak scenario is determined by the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ (in addition to eq. 7), which is not equivalent to a single requirement on the maximum sound speed $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$. For instance, a monotonically decreasing $C_{s}^{\mathrm{UD}}$ with $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$ does not lead to a second peak at high densities. In fact, monotonically decreasing $C_{s}$ between TOV and pQCD points are strongly disfavored, see table 7 in appendix D. The upshot is that a flat prior on $C_{s,\mathrm{max}}^{\mathrm{UD}}$ is not the least informative. Nevertheless, it is illuminating to explicitly exclude high sound speed by imposing upper bounds on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ in the prior. For the purpose of clarifying phase space volumes, let us demand $C_{s,\mathrm{max}}^{\mathrm{UD}}\leq 2/3$ by replacing the requirement on the prior eq. 14 with the following:
@@
The expression for $\Delta P_{\mathrm{min}}$ is related to eq. 11 via swapping subscripts TOV and pQCD and is given in ref [66]. The resulting Bayes factors are shown in table 4. Narrowing the phase space volume in the prior reduces the evidence, but the main findings largely remain unaffected.
@@
Table 5: Similar to table 1 but expands the correlated ${\chi\text{EFT}}$ uncertainty band from $2\sigma$ to $4\sigma$ when constructing NSs. The Bayes factors here are obtained from a (much) smaller pool of samples, and are largely similar to table 1.
@@
Table 6: Similar to table 1 but only follows ${\chi\text{EFT}}$ up to $1.5n_{0}$, above which density agnostic parameterizations  appendix B are used to describe NS inner cores. Even if ${\chi\text{EFT}}$ breaks down earlier, $M_{\text{TOV}}\leq 2.2M_{\odot}$ would still yield substantial evidence if first-order phase transitions are absent, see fig. 7. Constraints from GW170817 are approximated as a hard cut off $\Lambda_{1.4M_{\odot}}\leq 600$. This simplification renders the bound stronger than implied by the actual data, and is more unfavorable to the two-peak scenario. Yet the reduction to the evidence is mild, demonstrating that in agnostic scenarios the EOS at the TOV point largely decouples from that at lower densities relevant for most NSs [64, 66].
@@
The phase space for the two-peak scenario is also controlled by the lower bound of sound speed squared $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. So far, only $C_{s,\mathrm{min}}^{\mathrm{UD*}}=0$ required by thermodynamic stability is imposed. Low sound speed generally signals first-order phase transitions where low- and high-density phases in mechanical equilibrium are separated by a finite latent heat so that $C_{s}=\mathrm{d}P/\mathrm{d}\mathcal{E}\simeq 0$. Based on the similar symmetry breaking patterns and low-lying excitations in nucleon superfluid and in quark color superconductor, it has been conjectured that the quark-hadron transition at finite density is a smooth crossover [87, 88, 89] 555though topological phase transitions inside NSs are possible [113, 114, 115, 116], in which case the ultra-dense regime is known to be in the quark phase and likely free of additional phase transitions., justifying more restrictive $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. As discussed in the main text, when $C_{s,\mathrm{min}}^{\mathrm{UD*}}>0$ the area $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases rendering the criterion eq. 12 less demanding. The Bayes factors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ are shown in figs. 7 and 8. With $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.15$, even the unfavorable data $M_{\text{TOV}}\lesssim 2.2M_{\odot}$ would yield substantial support for two peaks. When $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, most EOSs are incompatible with $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.2$ [66], rendering the statistics less reliable with stronger assumptions about $C_{s,\mathrm{min}}^{\mathrm{UD*}}$.
@@
Other parameters of this framework are $n_{\chi\text{EFT}}$, the highest density ${\chi\text{EFT}}$ is adopted, and $\mu_{\text{pQCD}}$, the lowest chemical potential pQCD remains reliable. The convergence of ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=2n_{0}$ is demonstrated by statistical analyses exploiting the structure of EFT [85], although its uncertainty may have been underestimated [117]. The evidence is only marginally affected if the EFT error band is increased from $2\sigma$ to $4\sigma$, see table 6. Additionally, ${\chi\text{EFT}}$ may not be suitable if the ground state is no longer nucleonic at $\sim 2n_{0}$, such as the case of low-density quarkyonic transitions [40]. To examine these possibilities, I repeat the Bayesian analyses but only adopt ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=n_{0},1.5n_{0}$. The updated Bayes evidence is shown in table 6 and fig. 7. If the NS maximum mass is indeed low near $2.2M_{\odot}$, low-density nuclear theory inputs between $n_{0}$ and $2n_{0}$ would be helpful to secure clear evidence for two peaks in $C_{s}$.
@@
On the high-density end, $\mu_{\text{pQCD}}=2.4-2.6$ GeV is commonly adopted in the literature, although the large sensitivity on $X$ raises questions about this assumption. Refs  [55, 59] resummed to all order the leading and next-to-leading soft modes, and found reduced dependence on $X$ favoring $X\gtrsim 1.5$. Ref [34] inferred $X$ from comparing the N2LO pQCD EOS with lattice calculations at finite isospin [86, 33]. By including the LO pQCD prediction for the superconducting gap and exploiting the similar structure of pQCD at finite isospin and at finite $\mu_{B}$, $X$ is found to be large $X\in[2.9,4]$ and the superconducting gap at finite $\mu_{B}$ is found to be negligible with $\Delta\lesssim\text{MeV}$ near $\mu_{B}=2.4~{}\text{GeV}$. This bound on $X$ translates to a tightly constrained $P_{\text{pQCD}}\in[2.8,3.0]$ GeV/fm3 at $\mu_{B}=2.4~{}\text{GeV}$, which roughly corresponds to $X\gtrsim 2.6$ when soft modes at N3LO are included (the pQCD EOS adopted in the main text). According to table 1, such high values of $X$ lead to a strong preference for the two-peak scenario even with $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. On the other hand, the suppressed superconducting gap seems to suggest other mechanisms are behind the postulated second peak in $C_{s}$. If the superconducting gap is underestimated by LO pQCD, the inferred $X$ thus $P_{\text{pQCD}}$ is reduced when matching to lattice data at finite isopsin, but the full pressure $P_{\mathrm{CQM}}=P_{\text{pQCD}}+P_{\Delta}$ receives a pairing contribution $P_{\Delta}=\mu_{B}^{2}\Delta^{2}/(3\pi^{2})$ that is important if $\Delta\gtrsim 100$ MeV and $X\lesssim 1.5$. The bottom line is that pairing gaps much larger than inferred in ref [34] can still yield a large $\Delta P$ essential for the two-peak evidence, even though $X$ would be reduced. These large gaps are near the high end of current estimates, and are compatible with astrophysics [64, 65]. In short, while there exists strong indications supporting the reliability of pQCD down to $\mu_{B}=2.4$ GeV, higher order calculations and detailed statistical analyses will be essential to confirm the convergence of pQCD and to pinpoint the size of superconducting gaps.
@@
Figure 8 explores a pessimistic scenario and shows Bayes evidence if the matching to pQCD can only be performed at higher baryon chemical potentials $\mu_{B}=2.8,3.2$ GeV. The reduction of Bayes factors is considerable if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$, in which case only stronger assumptions $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.15-0.2$ may lead to substantial evidence. Once again, utilizing ab-initio calculations to the fullest extent is helpful to extract features in $C_{s}$ at ultra-high densities.
@@
Lastly, I demonstrate the robustness of the inferred bounds on the magnitude of the peak and trough. Figure 6 shows the posterior CIs if a flat prior on $\sqrt{C_{s,\mathrm{max}}^{\mathrm{UD}}}$ is imposed. Now, the mean of the prior is $C_{s,\mathrm{max}}^{\mathrm{UD*}}\approx 0.25$ favoring low sound speed. The posteriors are only marginally shifted, thus confirming the preference for high $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ as a physical consequence. More detailed investigations into the height of peaks and troughs in $C_{s}$ will be reported later.
Figure 7: Bayes factors based on neutral priors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. Each row assumes a different value of or distribution on $X$. By the geometric interpretation of $\Delta P$, higher $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ requires higher $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ to make up for the same area $\Delta P$. The evidence for two peaks thus becomes stronger. An exception is for $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, when thermodynamics demands $C_{s,\mathrm{min}}^{\mathrm{UD*}}\lesssim 0.1-0.2$ [66]. In this case significantly fewer samples exist so the evidence becomes more uncertain and may decrease with increasing $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. In the first column, both the mass and radius measurements of PSR J0740+6620 are imposed; the second and third columns are identical to the last two rows in table 1. Error bars reflect systematics associated with NS inner core models.
@@ Appendix D Illustrations of favored and disfavored scenarios
Here, I illustrate behaviors of $C_{s}$ at high densities that are favored and disfavored. The examples listed below are only demonstrations. Their functional form and numerical values are not to be construed as a posterior distribution or a bound, except that above NS densities the peak in $C_{s}$ exceeds $1/3$, or in scenarios with multiple peaks in the ultra-dense phase, at least one of them exceeds $1/3$. More refined constraints on $C_{s}(n_{B})$ will be reported in a subsequent paper.
Note that eq. 13 does not explicitly require $C_{s,\mathrm{max}}^{\mathrm{NS}}$ to be greater than $1/3$, though astrophysics strongly suggests the violation of conformal limit inside NSs. Moreover, while $C_{s}(n_{B})$ is allowed to have a finite number of discontinuities, for aesthetic reasons I only show smooth and continuous curves here. Finally, at low densities the possibility of a peak in $C_{s}$ in the ${\chi\text{EFT}}$-based EOS or in the crust is not counted towards the number of peaks in this work.

### v3

@@ B.5 Summary of models
This appendix examines assumptions underlie the main results in addition to those discussed in appendix B. I begin by discussing choices of prior distributions. In the main text, neutral priors without preferences for the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ are chosen. They are obtained by imposing separate flat distributions over regions where $\Delta P-\Delta P_{\mathrm{mean,max}}>0$ and where $\Delta P-\Delta P_{\mathrm{mean,max}}<0$. Normalizations of the two segments are obtained from demanding the integrated probability equals on both sides. The effect of eq. 7 is minor, and I have verified that neutral priors on $\Delta P-\Delta P_{\mathrm{mean,max}}$ alone lead to unbiased priors for the two-peak scenario.
Another common choice in the literature is a simple flat prior over possible ranges of $\Delta P-\Delta P_{\mathrm{mean,max}}$. It is deemed the least informative and generally preferred. Bayes factors based on this choice are shown in table 4. Here, the priors display clear preferences. The two-peak scenario is generally disfavored by these priors unless $X\gtrsim 3$. Although the statistical significance of the Bayesian evidence decreases mildly in these disadvantageous settings, the two-peak scenario is still favored.
@@
It is noteworthy that the criteria for the two-peak scenario eq. 13 are not equivalent to a single requirement on the maximum sound speed $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$. For instance, a monotonically decreasing $C_{s}^{\mathrm{UD}}$ with $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$ does not lead to a second peak at high densities. In fact, monotonically decreasing $C_{s}$ between TOV and pQCD points are strongly disfavored, see table 7 in appendix D. The upshot is that flat priors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ are not the least informative ones. Nevertheless, it is illuminating to explicitly exclude high sound speed by imposing upper bounds on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ in the prior. For the purpose of clarifying phase space volumes, let us demand $C_{s,\mathrm{max}}^{\mathrm{UD*}}\leq 2/3$ by replacing the requirement on the prior eq. 14 with the following:
@@
The expression for $\Delta P_{\mathrm{min}}$ is related to eq. 11 via swapping subscripts TOV and pQCD and is given in ref [66]. The resulting Bayes factors are shown in table 4. Narrowing the phase-space volume in the prior reduces the evidence, but the main findings remain unaffected.
@@
Table 5: Similar to table 1 but expands the correlated ${\chi\text{EFT}}$ uncertainty band from $2\sigma$ to $4\sigma$ when constructing NSs. Uniform distributions within the ${\chi\text{EFT}}$ uncertainty band are assumed. The Bayes factors here are obtained from a (much) smaller pool of samples, and are largely similar to table 1.
@@
Table 6: Similar to table 1 but only follows ${\chi\text{EFT}}$ up to $1.5n_{0}$, above which density agnostic parameterizations  appendix B are used to describe NS inner cores. The number of generated samples is roughly the same as when $n_{\chi\text{EFT}}=2n_{0}$ is assumed. Even if ${\chi\text{EFT}}$ breaks down earlier, $M_{\text{TOV}}\leq 2.2M_{\odot}$ would still yield substantial evidence if first-order phase transitions are absent, see fig. 7. Constraints from GW170817 are approximated as a hard cut off $\Lambda_{1.4M_{\odot}}\leq 600$. This simplification renders the bound stronger than implied by the actual data, and is more unfavorable to the two-peak scenario. Yet the reduction to the evidence is mild, demonstrating that with agnostic models the EOS at the TOV point largely decouples from that at lower densities relevant for most NSs [64, 66].
@@
The phase space for the two-peak scenario is also controlled by the lower bound of sound speed squared $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. So far, only $C_{s,\mathrm{min}}^{\mathrm{UD*}}=0$ required by thermodynamic stability is imposed. Low sound speed generally signals first-order phase transitions where low- and high-density phases in mechanical equilibrium are separated by a finite latent heat so that $C_{s}=\mathrm{d}P/\mathrm{d}\mathcal{E}\simeq 0$. Based on similar symmetry breaking patterns and low-lying excitations in nucleon superfluid and in quark color superconductor, it has been conjectured that the quark-hadron transition at finite density is a smooth crossover [87, 88, 89] 555though topological phase transitions are possible [114, 115, 116, 117]. The ultra-dense regime would still be free of phase transitions if they occur inside NSs., motivating more restrictive $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. As discussed in the main text, when $C_{s,\mathrm{min}}^{\mathrm{UD*}}>0$ the area $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases rendering the criterion eq. 12 less demanding. The Bayes factors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ are shown in figs. 7 and 8. With $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.15$, even the unfavorable data $M_{\text{TOV}}\lesssim 2.2M_{\odot}$ would yield substantial support for two peaks. When $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, most EOSs are incompatible with $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.2$ [66], rendering the statistics less reliable with stronger assumptions about $C_{s,\mathrm{min}}^{\mathrm{UD*}}$.
@@
Other parameters of this framework are $n_{\chi\text{EFT}}$, the highest density ${\chi\text{EFT}}$ is adopted, and $\mu_{\text{pQCD}}$, the lowest chemical potential pQCD remains reliable. The convergence of ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=2n_{0}$ is demonstrated by statistical analyses exploiting the structure of EFT [85], although its uncertainty may have been underestimated [118]. The evidence is only marginally affected if the EFT error band is broadened from $2\sigma$ to $4\sigma$, see table 6. Here and throughout this work, flat distributions within the ${\chi\text{EFT}}$ uncertainty band are assumed. This is achieved by imposing flat priors on the coefficients of basis functions in our ${\chi\text{EFT}}$-based low-density model, see eq 2 in [64]. Additionally, ${\chi\text{EFT}}$ may not be suitable if the ground state is no longer nucleonic at $\sim 2n_{0}$, such as the case of low-density quarkyonic transitions [40]. To examine these possibilities, I repeat the Bayesian analyses but only adopt ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=n_{0},1.5n_{0}$. The updated Bayes evidence is shown in table 6 and fig. 7. If the NS maximum mass is indeed low near $2.2M_{\odot}$, low-density nuclear theory inputs between $n_{0}$ and $2n_{0}$ would be helpful to secure clear evidence for two peaks in $C_{s}$.
@@
On the high-density end, $\mu_{\text{pQCD}}=2.4-2.6$ GeV is commonly adopted in the literature, although the large sensitivity on $X$ raises questions about this assumption. Refs  [55, 59] resummed to all order the leading and next-to-leading soft modes, and found reduced dependence on $X$ favoring $X\gtrsim 1.5$. Ref [34] inferred $X$ from comparing the N2LO pQCD EOS with lattice calculations at finite isospin [86, 33]. By including the LO pQCD prediction for the superconducting gap and exploiting the similar structure of pQCD at finite isospin and at finite $\mu_{B}$, $X$ is found to be large $X\in[2.4,4]$ and the superconducting gap at finite $\mu_{B}$ is found to be negligible with $\Delta\lesssim\text{MeV}$ near $\mu_{B}=2.4~{}\text{GeV}$. This bound on $X$ translates to a tightly constrained $P_{\text{pQCD}}\in[2.8,3.0]$ GeV/fm3 at $\mu_{B}=2.4~{}\text{GeV}$, which roughly corresponds to $X\gtrsim 2.3$ when soft modes at N3LO are included (the pQCD EOS adopted in the main text). According to table 1, such high values of $X$ lead to a notable preference for the two-peak scenario even with $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. On the other hand, the suppressed superconducting gap seems to suggest other mechanisms are behind the postulated second peak in $C_{s}$. If the superconducting gap is underestimated by LO pQCD, the inferred $X$ thus $P_{\text{pQCD}}$ is reduced when matching to lattice data at finite isopsin, but the full pressure $P_{\mathrm{CQM}}=P_{\text{pQCD}}+P_{\Delta}$ receives a pairing contribution $P_{\Delta}=\mu_{B}^{2}\Delta^{2}/(3\pi^{2})$ that is important if $\Delta\gtrsim 100$ MeV and $X\lesssim 1.5$. The bottom line is that pairing gaps much larger than inferred in ref [34] can still yield a large $\Delta P$ essential for the two-peak evidence, even though $X$ would be reduced. These large gaps are near the high end of current estimates, and are compatible with astrophysics [64, 65]. In short, while there exists strong indications supporting the reliability of pQCD down to $\mu_{B}=2.4$ GeV, higher order calculations and detailed statistical analyses will be essential to assess the applicable range of asymptotic expansions in $\alpha_{s}$ and to pinpoint the size of superconducting gaps.
@@
Figure 8 explores a pessimistic scenario and shows Bayes evidence if the matching to pQCD can only be performed at higher baryon chemical potentials $\mu_{B}=2.8,3.2$ GeV. The reduction of Bayes factors is considerable if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$, in which case only stronger assumptions $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.15-0.2$ can lead to substantial evidence. Once again, utilizing ab-initio calculations to the fullest extent is helpful to extract features in $C_{s}$ at ultra-high densities.
@@
Lastly, I demonstrate the robustness of the inferred bounds on the magnitude of the peak and trough. Figure 6 shows the posterior CIs if a flat prior on $\sqrt{C_{s,\mathrm{max}}^{\mathrm{UD}}}$ is imposed. Now, the mean of the prior is $C_{s,\mathrm{max}}^{\mathrm{UD*}}\approx 0.25$ favoring low sound speed. The posteriors are only marginally shifted, confirming that the preference for high $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ is a physical consequence. More detailed investigations into the height of peaks and troughs in $C_{s}$ will be reported later.
Figure 7: Bayes factors based on neutral priors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. Each row assumes a different value of or distribution on $X$. By the geometric interpretation of $\Delta P$, higher $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ requires higher $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ to make up for the same area $\Delta P$. The evidence for two peaks thus becomes stronger. An exception is for $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, when thermodynamics demands $C_{s,\mathrm{min}}^{\mathrm{UD*}}\lesssim 0.1-0.2$ [66]. In this case, there are significantly fewer samples, so the evidence becomes more uncertain and may decrease with increasing $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. In the first column, both the mass and radius measurements of PSR J0740+6620 are imposed; the second and third columns are identical to the last two rows in table 1. Error bars reflect systematics associated with NS inner core models.
@@ Appendix D Illustrations of favored and disfavored scenarios
Here, I illustrate behaviors of $C_{s}$ at high densities that are favored and disfavored. The examples listed below are only demonstrations. Their functional form and numerical values are not to be construed as posterior distributions or bounds. For instance, the value of $n_{\text{TOV}}$ is model-dependent and in reality would not be identical for the examples shown below. The only requirement is the existence of a peak in $C_{s}$ exceeding $1/3$ following a trough above $n_{\text{TOV}}$. More refined constraints on $C_{s}(n_{B})$ will be reported in a subsequent paper.
Note that eq. 13 does not explicitly require $C_{s,\mathrm{max}}^{\mathrm{NS}}$ to be greater than $1/3$, though astrophysics strongly suggests the violation of conformal limit inside NSs. Moreover, while $C_{s}(n_{B})$ is allowed to have a finite number of discontinuities, a fact implicit in the central results  eq. 13 and in NS models in  appendix B (via very short length scales $\tilde{l}$), for aesthetic reasons I only show smooth curves here. Finally, at low densities the possibility of a peak in $C_{s}$ in the ${\chi\text{EFT}}$-based outer core or in the crust is not counted towards the number of peaks in this work.